# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ C.3 Magnetic field configuration
At very large radial distance on the disc, the emitter can be approximated as static, so that $\chi=0$ and $\omega=-\boldsymbol{k}\cdot\boldsymbol{u}=-g_{tt}k^{t}$. Also, since far from the black hole the spacetime is described by the Minkowski metric, with $g_{tt}=-1$ and $g_{rr}=1$, we get $\omega=k^{t}$ and $k^{t}$ can be related to the photon energy defined in equation (2.1): $k^{t}=g^{tt}k_{t}=-E/(g_{tt}\hbar)=E/\hbar$. Using all these remarks, equations (C2) drastically simplify in the asymptotic limit $r\to\infty$: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Vertical}}=-\hbar rk^{\theta}/E$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Radial}}=\hbar k^{r}/E$. Then, noticing that for a face-on observer the angular momentum $L$ of equation (2.1) is zero for all photons, one can retrieve the expressions of $k^{\theta}$ and $k^{r}$ by taking the vanishing spin parameter $a$ in the Kerr equations of motion (Gralla & Lupsasca 2020b). We finally get: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Vertical}}=\mp_{\theta}\sqrt{\bar{Q}}/r$, with the minus sign corresponding to a photon emitted downwards in $\theta$, and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Radial}}=\pm_{r}\sqrt{1-\bar{Q}(r^{2}-2Mr)/r^{4}}$, with the minus sign corresponding to a photon emitted towards the black hole radially. For an on-axis observer, the reduced Carter constant $\bar{Q}$ is directly related to the polar radius $b$ on the screen (Gralla & Lupsasca 2020a) via $\bar{Q}=b^{2}$ and the approximation (2.3), $b\approx r+M$, then yields: $\bar{Q}\approx r^{2}+2Mr+M^{2}$. Thus, for $n=0$ photons, $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=0}_{\text{Vertical}}=-1\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=0}_{\text{Vertical}}=0$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=0}_{\text{Radial}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=0}_{\text{Radial}}=1$. For $n=1$ photons, it is clear from Figure C.3 that the radial growth of $\bar{Q}$, unchanged along one geodesic, is much slower than $r^{2}$ with respect to the second impact radius. As a consequence, $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=1}_{\text{Vertical}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=1}_{\text{Vertical}}=1$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=1}_{\text{Radial}}=-1\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=1}_{\text{Radial}}=0$, where the ingoing sign can be understood from Figure C.3. When radially approaching the event horizon, no trivial [^34] Except for the conditions $u^{\theta}=0$ and $L=p_{\varphi}=g_{\varphi\varphi}p^{\varphi}=0$, translating an equatorial disc motion and a face-on view respectively. simplifications apply to $\omega=-g_{tt}k^{t}u^{t}-g_{rr}k^{r}u^{r}$. However, again from the geodesic equations (Gralla & Lupsasca 2020b), the asymptotic behaviour of $w$ is $w^{r_{\mathrm{H}}}=[u^{t}-u^{r}r/(r-2M)]E/\hbar$. Then, using the fact that $u^{t}>0$ for future-directed particles and $u^{r}$ is negative and non-vanishing below the radius of the ISCO, it is straightforward to conclude that $\omega\to+\infty$ at the horizon. As $k^{\theta}$ is finite at the horizon, for a vertical magnetic field, we find: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Vertical}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Vertical}}=1$. In the case of a radial magnetic field, we show that the asymptotic limit of equation (C2) is $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Radial}}=\varepsilon_{\mathrm{sub}}r\left(\varepsilon_{\mathrm{sub}}^{2}r^{2}+C^{2}\omega_{\varphi}^{2}\ell_{\mathrm{sub}}^{2}\right)^{-1/2}$ where we have introduced $C\coloneqq-\omega_{r}\varepsilon_{\mathrm{sub}}-(1-\omega_{r})$ and employed the quantities $\varepsilon_{\mathrm{sub}},\ell_{\mathrm{sub}},\omega_{r}$ and $\omega_{\varphi}$ presented in Paragraph 4.1.3. For a sub-Keplerian motion, for which $\omega_{r}=\omega_{\varphi}=1$, this quantity reduces to $r\left(r^{2}+\ell_{\mathrm{sub}}^{2}\right)^{-1/2}=\frac{1}{2}[(4-\xi^{2})/(1+2\xi^{2})]^{1/2}$ which is equal to 0.5 for Keplerian velocity profile with $\xi=1$, see Paragraph 4.1.3. For a radial infall, the previous scalar product becomes 1 as $\omega_{r}=\omega_{\varphi}=0$. To summarise, given the values of Table 2, $(\sin\theta_{\mathrm{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Radial}}\simeq 0.87$, 0 or 0.67 for a Keplerian, radial or mixed velocity profile respectively. Let us finally consider the intermediate extrema of the $\sin\theta_{\mathrm{B}}$ function. Generally, an extremum of $\sin\theta_{\mathrm{B}}$ is found either when $\cos\theta_{\mathrm{B}}=0\Leftrightarrow\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}}=0$ or when $\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}}\neq\pm 1$ and $d(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})/dr=0$. The first condition, corresponding to a maximum as $\sin\theta_{\mathrm{B}}=1$, is never verified for a vertical magnetic field, as $\omega$ is finite and $k^{\theta}\neq 0$ for non-extremal values of $r$, and it is satisfied, in the case of a radial magnetic field for $k^{r}/k^{t}=\chi=u^{r}/u^{t}$. We numerically show that, for all the velocity profiles investigated in this paper, this relation is never verified for the $n=0$ photons and only for a single value of $r$, depending on the disc velocity, in the case of $n=1$ photons. In particular, $r\simeq 3.6$, 5.6 or $4.4M$ in the case, respectively, of a Keplerian, radial or mixed velocity with the parameters of Table 2. In the case of a vertical magnetic field, using Leibniz rule and the geodesic equations (Gralla & Lupsasca 2020b), the second condition can be rewritten as $r=-\omega/(d\omega/dr)$. We note that this condition is never verified for $n=0$ photons, whereas it leads to a single [^35] As $(\sin\theta_{\mathrm{B}})^{n=1}_{\text{Vertical}}$ is not a constant function and it tends to 1 both at infinity and at the horizon, we already knew that it possesses at least one minimum; here, we claim the unicity of this extremum. minimum position $r>2M$ for $n=1$ photons: $r\simeq 3.9$, 5.3 or $4.6M$ for a Keplerian, radial or mixed velocity, respectively. For a radial magnetic field, no trivial simplification of the second condition can be easily found, but we notice that an additional minimum is present for the Keplerian and mixed velocity profiles both for the $n=0$ and $n=1$ photons at, respectively, $r\simeq 2.6M$ and $r\simeq 2.7M$ or $r\simeq 2.2M$ and $r\simeq 2.3M$. Let us now summarise our findings and their implications for the position of the peaks of intensity. For weakly lensed photons and a radial magnetic field, $\sin\theta_{\mathrm{B}}$ increases from an intermediate value to 1 with increasing distance from the central object, after reaching a minimum in the case of a Keplerian or mixed velocity, while it decreases from 1 to 0 in the presence of a vertical magnetic field. As a consequence, $j_{\nu}$ given by equation (21), at a fixed $r$, is smaller, in both cases, than when taking $\theta_{\mathrm{B}}=90^{\circ}$, but more importantly, the emission profile dies faster for a radial magnetic field and even more so for a vertical one. This implies that, when incorporating redshift effects, the $n=0$ peak positions will be reached at further radii for the $\theta_{\mathrm{B}}=90^{\circ}$ profiles than for the ones where $\theta_{\mathrm{B}}$ is computed numerically. Specifically, the maxima occur at smaller radii in the case of a radial magnetic field and even smaller radii for a vertical field. Part of this reasoning is confirmed by comparing the left columns of Figures 7 and 8. Concerning the first-lensed photons, whose geodesics are illustrated in light blue in Figure C.3, a global extremum in $\sin\theta_{\mathrm{B}}$ is present in all scenarios, being a minimum for a vertical magnetic field and a maximum for a radial one, see right inset of Figure C.3 for a mixed velocity profile. Then, in the radial-$B$ case, inner emission below the maximum is weakened even before the redshift knock down detailed in Section 6. This explains the behaviour remarked in Figure C.2: $n=1$ peak positions appear further away for a radial magnetic field. Similarly, as the inner emission for a vertical magnetic field is boosted, $n=1$ peaks appear at smaller radii than the ones of profiles with constant $\theta_{\mathrm{B}}=90^{\circ}$; see right columns of Figures 7 and 8. The integration of $\theta_{\mathrm{B}}$ in the thermal synchrotron emission profile is then an essential ingredient to take into account when interpreting observed images and its imprints on lensed observables could provide some hints on the magnetic field structure.
@@ C.4 $r_{\mathrm{inner}}$
As already pointed out in Section 7.1, the key ingredient to break the degeneracy with the peak positions of the first photon ring is the different size of the critical curve in various spacetimes. The question now is to know how robust this result is when varying the position of the inner radius of the disc. For example, by taking a RZ spacetime with $\epsilon\,|_{\,a_{1}=0}=0.1$ and a disc with an inner radius $r_{\mathrm{inner}}=2M$ instead of $1.81M$, coinciding with the horizon of a Schwarzschild black hole, the violet band of the bottom right panel of Figure 8 is shifted upwards by at most 0.05 µas, preserving the conclusions about possible observable disentanglements. However, already for $r_{\mathrm{inner}}=3M$, the violet band is just 0.1 µas below the pink one, meaning that the impacts of $a_{1}$ and $\epsilon$ become comparable. The position of the inner radius of the disc is even more crucial for the analysis of the primary image: unlike variations in the critical curve, which do not impact the $n=0$ peak positions, the latter are primarily governed by the inner radius, where the considered emission profiles reach their maximum intensity. For instance, in the case, again, of a RZ spacetime with $\epsilon\,|_{\,a_{1}=0}=0.1$ and a disc with $r_{\mathrm{inner}}=2M$, the the violet band of the bottom left panel of Figure 8 gets closer to the pink one with a shift of about 0.25 µas, putting the deviations on $a_{1}$ and $\epsilon$ on a more similar footing. For $r_{\mathrm{inner}}=3M$, the upper part of the Schwarzschild green band is encompassed. Hence, the position of the inner edge of the accretion disc could be an important source of uncertainty, especially for the interpretation of the direct image. Nevertheless, for MAD discs, the presence of dynamically important near-horizon magnetic fields ensures that the plasma moves slowly all the way down to the BH horizon (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2021), thus taking $r_{\mathrm{inner}}=r_{\mathrm{H}}$ is a fair choice.
@@ C.5 Inclination
We analyse here the contribution of a strong inclination, namely $i=100^{\circ}$, to the study of the degeneracy on the peaks of the 1D intensity profiles. Although this viewing angle is not applicable to M87* or Sgr A* (The Event Horizon Telescope Collaboration 2022), the investigations of this paragraph could be useful when, with the future interferometric BHEX observations, a broader number of supermassive black holes will be within reach (Johnson et al. 2024). To properly account for the deformation due to projection effects, extensively discussed in Appendix A, the edges of the first lensing band are ray-traced on 100 polar angles on the screen instead of 14 as in Section 5; on the other hand, the radial resolution is unchanged. It is interesting to notice that, contrary to the low-inclination framework treated in the main text, here there are some privileged polar directions on the screen to look for metric deviations. In Figure C.4, the three bands are clearly separated on the left side, while they are widely superposed on the right. As it is quite evident from Figure C.4, the fractional asymmetry is higher than for $i=163^{\circ}$, reaching almost 40% for the primary image and 3% for the first photon ring. Those values are nevertheless degenerate between the various spacetimes as well as those of the size ratios.
Figure 16: Positions of the $n=1$ peaks, observed at $i=100^{\circ}$, in the case of a Schwarzschild BH (green), a RZ BH with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ BH with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Critical curves are dashed lines.

### v2

@@ C.3 Magnetic field configuration
At very large radial distance on the disc, the emitter can be approximated as static, so that $\chi=0$ and $\omega=-\boldsymbol{k}\cdot\boldsymbol{u}=-g_{tt}k^{t}$. Also, since far from the black hole the space-time is described by the Minkowski metric, with $g_{tt}=-1$ and $g_{rr}=1$, we got $\omega=k^{t}$ and $k^{t}$ could be related to the photon energy defined in equation (2.1): $k^{t}=g^{tt}k_{t}=-E/(g_{tt}\hbar)=E/\hbar$. Using all these remarks, equations (C2) drastically simplify in the asymptotic limit $r\to\infty$: $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Vertical}}=-\hbar rk^{\theta}/E$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Radial}}=\hbar k^{r}/E$. Then, noticing that for a face-on observer the angular momentum, $L$, of equation (2.1) is zero for all photons, we could retrieve the expressions of $k^{\theta}$ and $k^{r}$ by taking the vanishing spin parameter, $a$, in the Kerr equations of motion (Gralla & Lupsasca 2020b). We finally got $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Vertical}}=\mp_{\theta}\sqrt{\bar{Q}}/r$, with the minus sign corresponding to a photon emitted downwards in $\theta$, and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty}_{\text{Radial}}=\pm_{r}\sqrt{1-\bar{Q}(r^{2}-2Mr)/r^{4}}$, with the minus sign corresponding to a photon emitted towards the black hole radially. For an on-axis observer, the reduced Carter constant, $\bar{Q}$, is directly related to the polar radius, $b$, on the screen (Gralla & Lupsasca 2020a) via $\bar{Q}=b^{2}$ and the approximation (2.3), $b\approx r+M$, then yields $\bar{Q}\approx r^{2}+2Mr+M^{2}$. Thus, for $n=0$ photons, $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=0}_{\text{Vertical}}=-1\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=0}_{\text{Vertical}}=0$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=0}_{\text{Radial}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=0}_{\text{Radial}}=1$. For $n=1$ photons, it is clear from Figure C.3 that the radial growth of $\bar{Q}$, unchanged along one geodesic, is much slower than $r^{2}$ with respect to the second impact radius. As a consequence, $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=1}_{\text{Vertical}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=1}_{\text{Vertical}}=1$ and $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{\infty,\,n=1}_{\text{Radial}}=-1\Rightarrow(\sin\theta_{\mathrm{B}})^{\infty,\,n=1}_{\text{Radial}}=0$, where the ingoing sign can be understood from Figure C.3. When radially approaching the event horizon, no trivial [^34] Except for the conditions $u^{\theta}=0$ and $L=p_{\varphi}=g_{\varphi\varphi}p^{\varphi}=0$, translating an equatorial disc motion and a face-on view respectively. simplifications apply to $\omega=-g_{tt}k^{t}u^{t}-g_{rr}k^{r}u^{r}$. However, again from the geodesic equations (Gralla & Lupsasca 2020b), the asymptotic behaviour of $w$ is $w^{r_{\mathrm{H}}}=[u^{t}-u^{r}r/(r-2M)]E/\hbar$. Then, using the fact that $u^{t}>0$ for future-directed particles and $u^{r}$ is negative and non-vanishing below the radius of the ISCO, it is straightforward to conclude that $\omega\to+\infty$ at the horizon. As $k^{\theta}$ is finite at the horizon, for a vertical magnetic field, we find $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Vertical}}=0\Rightarrow(\sin\theta_{\mathrm{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Vertical}}=1$. In the case of a radial magnetic field, we show that the asymptotic limit of equation (C2) is $(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Radial}}=\varepsilon_{\mathrm{sub}}r\left(\varepsilon_{\mathrm{sub}}^{2}r^{2}+C^{2}\omega_{\varphi}^{2}\ell_{\mathrm{sub}}^{2}\right)^{-1/2}$ where we have introduced $C\coloneqq-\omega_{r}\varepsilon_{\mathrm{sub}}-(1-\omega_{r})$ and employed the quantities $\varepsilon_{\mathrm{sub}},\ell_{\mathrm{sub}},\omega_{r}$, and $\omega_{\varphi}$ presented in Paragraph 4.1.3. For a sub-Keplerian motion, for which $\omega_{r}=\omega_{\varphi}=1$, this quantity reduces to $r\left(r^{2}+\ell_{\mathrm{sub}}^{2}\right)^{-1/2}=\frac{1}{2}[(4-\xi^{2})/(1+2\xi^{2})]^{1/2}$ which is equal to 0.5 for Keplerian velocity profile with $\xi=1$ (see Paragraph 4.1.3). For a radial infall, the previous scalar product becomes 1 as $\omega_{r}=\omega_{\varphi}=0$. To summarise, given the values of Table 2, $(\sin\theta_{\mathrm{B}})^{r_{\mathrm{H}},\,n=0,1}_{\text{Radial}}\simeq 0.87$, 0 or 0.67 for a Keplerian, radial or mixed velocity profile respectively. Finally, we considered the intermediate extrema of the $\sin\theta_{\mathrm{B}}$ function. Generally, an extremum of $\sin\theta_{\mathrm{B}}$ is found when $\cos\theta_{\mathrm{B}}=0\Leftrightarrow\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}}=0$ or when $\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}}\neq\pm 1$ and $d(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})/dr=0$. The first condition, corresponding to a maximum as $\sin\theta_{\mathrm{B}}=1$, is never verified for a vertical magnetic field, as $\omega$ is finite and $k^{\theta}\neq 0$ for non-extremal values of $r$, and it is satisfied, in the case of a radial magnetic field for $k^{r}/k^{t}=\chi=u^{r}/u^{t}$. We numerically show that, for all the velocity profiles investigated in this paper, this relation is never verified for the $n=0$ photons and only for a single value of $r$, depending on the disc velocity, in the case of $n=1$ photons. In particular, $r\simeq 3.6$, 5.6 or $4.4M$ in the case, respectively, of a Keplerian, radial or mixed velocity with the parameters of Table 2. In the case of a vertical magnetic field, using Leibniz rule and the geodesic equations (Gralla & Lupsasca 2020b), the second condition can be rewritten as $r=-\omega/(d\omega/dr)$. We note that this condition is never verified for $n=0$ photons, whereas it leads to a single [^35] As $(\sin\theta_{\mathrm{B}})^{n=1}_{\text{Vertical}}$ is not a constant function and it tends to 1 both at infinity and at the horizon, we already knew that it possesses at least one minimum; here, we claim the unicity of this extremum. minimum position $r>2M$ for $n=1$ photons: $r\simeq 3.9$, 5.3 or $4.6M$ for a Keplerian, radial or mixed velocity, respectively. For a radial magnetic field, no trivial simplification of the second condition can be easily found, but we notice that an additional minimum is present for the Keplerian and mixed velocity profiles both for the $n=0$ and $n=1$ photons at, respectively, $r\simeq 2.6M$ and $r\simeq 2.7M$ or $r\simeq 2.2M$ and $r\simeq 2.3M$. Here, we summarise our findings and their implications for the position of the peaks of intensity. For weakly lensed photons and a radial magnetic field, $\sin\theta_{\mathrm{B}}$ increases from an intermediate value to 1 with increasing distance from the central object, after reaching a minimum in the case of a Keplerian or mixed velocity, while it decreases from 1 to 0 in the presence of a vertical magnetic field. As a consequence, $j_{\nu}$ given by equation (21), at a fixed $r$, is smaller, in both cases, than when taking $\theta_{\mathrm{B}}=90^{\circ}$, but more importantly, the emission profile dies faster for a radial magnetic field and even more so for a vertical one. This implies that, when incorporating redshift effects, the $n=0$ peak positions will be reached at further radii for the $\theta_{\mathrm{B}}=90^{\circ}$ profiles than for the ones where $\theta_{\mathrm{B}}$ is computed numerically. Specifically, the maxima occur at smaller radii in the case of a radial magnetic field and even smaller radii for a vertical field. Part of this reasoning is confirmed by comparing the left columns of Figures 7 and 8. Concerning the first-lensed photons, whose geodesics are illustrated in light blue in Figure C.3, a global extremum in $\sin\theta_{\mathrm{B}}$ is present in all scenarios, being a minimum for a vertical magnetic field and a maximum for a radial one (see right inset of Figure C.3 for a mixed velocity profile). Then, in the radial-$B$ case, inner emission below the maximum is weakened even before the redshift knock down detailed in Section 6. This explains the behaviour remarked in Figure C.2: $n=1$ peak positions appear further away for a radial magnetic field. Similarly, as the inner emission for a vertical magnetic field is boosted, $n=1$ peaks appear at smaller radii than the ones of profiles with constant $\theta_{\mathrm{B}}=90^{\circ}$ (see right columns of Figures 7 and 8). The integration of $\theta_{\mathrm{B}}$ in the thermal synchrotron emission profile is then an essential ingredient to take into account when interpreting observed images and its imprints on lensed observables could provide some hints on the magnetic field structure.
@@ C.4 $r_{\mathrm{inner}}$
As already pointed out in Section 7.1, the key ingredient to break the degeneracy with the peak positions of the first photon ring is the different size of the critical curve in various space-times. The question now is to know how robust this result is when varying the position of the inner radius of the disc. For example, by taking a RZ space-time with $\epsilon\,|_{\,a_{1}=0}=0.1$ and a disc with an inner radius $r_{\mathrm{inner}}=2M$ instead of $1.81M$, coinciding with the horizon of a Schwarzschild black hole, the violet band of the bottom right panel of Figure 8 is shifted upwards by at most 0.05 µas, preserving the conclusions about possible observable disentanglements. However, already for $r_{\mathrm{inner}}=3M$, the violet band is just 0.1 µas below the pink one, meaning that the impacts of $a_{1}$ and $\epsilon$ become comparable. The position of the inner radius of the disc is even more crucial for the analysis of the primary image. While variations in the critical curve do not impact the $n=0$ peak positions, they are primarily governed by the inner radius, where the considered emission profiles reach their maximum intensity. For instance, in the case, again, of a RZ space-time with $\epsilon\,|_{\,a_{1}=0}=0.1$ and a disc with $r_{\mathrm{inner}}=2M$, the violet band of the bottom left panel of Figure 8 gets closer to the pink one with a shift of about 0.25 µas, putting the deviations on $a_{1}$ and $\epsilon$ on a more similar footing. For $r_{\mathrm{inner}}=3M$, the upper part of the Schwarzschild green band is encompassed. Hence, the position of the inner edge of the accretion disc could be an important source of uncertainty, especially for the interpretation of the direct image. Nevertheless, for MAD discs, the presence of dynamically important near-horizon magnetic fields ensures that the plasma moves slowly all the way down to the black hole horizon (Narayan et al. 2003; The Event Horizon Telescope Collaboration 2021), thus taking $r_{\mathrm{inner}}=r_{\mathrm{H}}$ is a fair choice.
@@ C.5 Inclination
Figure 16: Positions of the $n=1$ peaks, observed at $i=100^{\circ}$, for all possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink), and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Critical curves are dashed lines.
We analysed the contribution of a strong inclination, namely $i=100^{\circ}$, to the study of the degeneracy on the peaks of the 1D intensity profiles. Although this viewing angle is not applicable to M87* or Sgr A* (The Event Horizon Telescope Collaboration 2022), the investigations of this paragraph could be useful when, with the future interferometric BHEX observations, a broader number of supermassive black holes will be within reach (Johnson et al. 2024). To properly account for the deformation due to projection effects, extensively discussed in Appendix A, the edges of the first lensing band were ray-traced on 100 polar angles on the screen instead of 14 as in Section 5; on the other hand, the radial resolution was unchanged. It is interesting to notice that, contrary to the low-inclination framework treated in the main text, here there are some privileged polar directions on the screen to look for metric deviations. In Figure C.4, the three bands are clearly separated on the left side, while they are widely superposed on the right. As it is quite evident from Figure C.4, the fractional asymmetry is higher than for $i=163^{\circ}$, reaching almost 40% for the primary image and 3% for the first photon ring. Those values are nevertheless degenerate between the various space-times as well as those of the size ratios.